# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 3 Hyperfocal abelian Frobenius blocks
Let $G$ be a finite group and $b$ a block of $G$ over $\mathcal{O}$ with defect pointed group $P_{\gamma}$. We denote the hyperfocal subgroup of $P_{\gamma}$ by $D$. Then by [25, Propositions 4.2 and 3.3], there is a unique local point $\delta$ of $D$ on $\mbox{$\mathcal{O}$}Gb$ such that $D_{\delta}\leq P_{\gamma}$. Take an element $i$ of $\gamma$ and set $A=i\mbox{$\mathcal{O}$}Gi$. Then $A$ is a source algebra of $\mbox{$\mathcal{O}$}Gb$ and let $\mathbb{A}$ be a hyperfocal subalgebra of $A$. We take an element $j$ of $\delta$ such that $ij=j=ji$. By [25, Proposition 2.8], we can assume that $j$ belongs to $\mathbb{A}$. We can view $D_{\delta}$ as a local pointed group of $D$ on both $A$ and $\mathbb{A}$. Set $A_{\delta}=jAj$ and $\mathbb{A}_{\delta}=j\mathbb{A}j$. Clearly, $A_{\delta}$ and $\mathbb{A}_{\delta}$ are embedded algebras of the local pointed group $D_{\delta}$ on $A$ and $\mathbb{A}$, respectively. By the uniqueness of $D_{\delta}$, there is an element $a_{u}$ in $(\mathbb{A}^{D})^{*}$ such that $a_{u}ja_{u}^{-1}=uju^{-1}$ for any $u\in P$. Denote $a_{u}^{-1}uj$ by $a_{u,\delta}$ for any $u\in P$. Then $a_{u,\delta}$ belongs to $A_{\delta}^{*}$ fulfilling $a_{u,\delta}(vj)a_{u,\delta}^{-1}=(uvu^{-1})j$ for any $v\in D$. Moreover, we have $A_{\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}_{\delta}a_{u,\delta}$ as $\mathcal{O}$-modules.
@@
It is well-known that the source algebra $A$ is closely related to the block algebra $\mbox{$\mathcal{O}$}Gb$. For example, they are both relatively $\mbox{$\mathcal{O}$}P$-separable $P$-interior algebras; there is a canonical Morita equivalence between the source algebra $A$ and the block algebra $\mbox{$\mathcal{O}$}Gb$ induced by the $A$-$\mbox{$\mathcal{O}$}Gb$-bimodule $i\mbox{$\mathcal{O}$}G$. Next, we will show that these properties also hold for the hyperfocal subalgebra $\mathbb{A}$ and its embedded subalgebra $\mathbb{A}_{\delta}$ by the following general result.
@@
Let $Q$ be a $p$-subgroup and $C$ an indecomposable relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebra. Assume that $C$ has a unique local pointed group $Q_{\epsilon}$ and as $\mbox{$\mathcal{O}$}(Q\times Q)$-modules, the indecomposable direct summand of $C$ has the form $\mathrm{Ind}_{\Delta_{\varphi}(T)}^{Q\times Q}(\mbox{$\mathcal{O}$})$ for some subgroup $T$ of $Q$ and some injective homomorphism $\varphi:T\longrightarrow Q$. Here, $\Delta_{\varphi}(T)=\{(u,\varphi(u))\mid u\in T\}$. Moreover, suppose that $\frac{\mathrm{rank}_{\mathcal{O}}(C)}{|Q|}$ is coprime to $p$. Take an element $l$ of $\epsilon$. Then the $Q$-interior algebra $lCl$ is a relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebra and the lCllCl-$C$-bimodule lClC induces a Morita equivalence between $lCl$ and $C$.
@@
Since the indecomposable direct summand of the $\mbox{$\mathcal{O}$}(Q\times Q)$-module $C$ has the form $\mathrm{Ind}_{\Delta_{\varphi}(T)}^{Q\times Q}(\mbox{$\mathcal{O}$})$, $C$ is a direct sum of some copies of $\mbox{$\mathcal{O}$}Q$ as left $\mbox{$\mathcal{O}$}Q$ and right $\mbox{$\mathcal{O}$}Q$-modules. In particular, $C$ is free as left and right $\mbox{$\mathcal{O}$}Q$-modules. Hence, the structural map $Q\rightarrow C^{*}$ is injective and we can view $\mbox{$\mathcal{O}$}Q$ as a subalgebra of $C$. Obviously, every finitely generated projective $C$-module has $\mathcal{O}$-rank divided by $|Q|$.
@@
By the definition of relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebras, $C$ is a direct summand of $C\mathop{\otimes}\limits_{\mathcal{O}Q}C$ as $C$-$C$-bimodules. Since $C$ is an indecomposable algebra, there are some primitive idempotents $e,f$ of $C^{Q}$ such that $C$ is a direct summand of $Ce\mathop{\otimes}\limits_{\mathcal{O}Q}fC$. Suppose that $\mathrm{Br}_{Q}(e)=0$. By [29, Theorem (23.1)], there is a proper subgroup $Q_{0}$ of $Q$ and a primitive idempotent $e_{0}$ of $C^{Q_{0}}$ such that $e=\mathrm{Tr}_{Q_{0}}^{Q}(j_{0})$ and ${{}^{u}}e_{0}\cdot e=0$ for any $u\in Q-Q_{0}$. Then the map sending $c\in Ce$ to $\sum\limits_{u\in Q/Q_{0}}cue_{0}\mathop{\otimes}\limits_{\mathcal{O}Q_{0}}u^{-1}$ induces a $C$-$\mbox{$\mathcal{O}$}Q$-bimodule isomorphism from $C_{e}$ to $Ce_{0}\mathop{\otimes}\limits_{\mathcal{O}Q_{0}}\mbox{$\mathcal{O}$}Q$. Therefore, $C$ is a direct summand of $Ce_{0}\mathop{\otimes}\limits_{\mathcal{O}Q_{0}}fC$ as $C$-$C$-bimodules. Denote by $\bar{C}$ the $k$-algebra $C/J(\mbox{$\mathcal{O}$})C$. Then for any simple $C$-module $W$, we have $W$ is a direct summand of $\bar{C}\mathop{\otimes}\limits_{kQ_{0}}U$ for some indecomposable $kQ_{0}$-module $U$. By assumption and the Mackey decomposition formula, the indecomposable direct summand of $\mathrm{Res}_{Q\times Q_{0}}^{Q\times Q}(\bar{C})$ has the form $kQ\mathop{\otimes}\limits_{kT}({i}kQ_{0})$ for some proper subgroup $T$ of $Q$ and injective homomorphism $\psi:T\longrightarrow Q_{0}$. By Green’s indecomposability theorem, this implies that the dimension of every simple $C$-module has dimension divided by $p$. Let $P_{W}$ be the projective cover of the simple $C$-module $W$. We have $\mathrm{dim}(\bar{C})=\sum\limits_{W}\mathrm{dim}(P_{W})\mathrm{dim}(W)$ which is divided by $p|Q|$. This is contradict the assumption that $\frac{\mathrm{rank}_{\mathcal{O}}(C)}{|Q|}$ is coprime to $p$.
@@
We conclude that $\mathrm{Br}_{Q}(e)\neq 0$. Similarly, we can also get $\mathrm{Br}_{Q}(f)\neq 0$. Since $\epsilon$ is the unique local point of $Q$ on $C$, we have $e$ and $f$ both belongs to $\epsilon$. In particular, we have $Ce$ is isomorphic to $Cl$ as $C$-$\mbox{$\mathcal{O}$}Q$-bimodules and $fC$ is isomorphic to $lC$ as $\mbox{$\mathcal{O}$}Q$-$C$-bimodules. So $C$ is a direct summand of $Cl\mathop{\otimes}\limits_{\mathcal{O}Q}lC$ as $C$-$C$-bimodules and then $lCl$ is a direct summand of $lCl\mathop{\otimes}\limits_{\mathcal{O}Q}lCl$ as lCllCl-lCllCl-bimodules. This means that $lCl$ is a relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebra.
@@
Since $C$ is a direct summand of $Cl\mathop{\otimes}\limits_{\mathcal{O}Q}lC$, $W$ is a direct summand of $Cl\mathop{\otimes}\limits_{\mathcal{O}Q}l\cdot W$ for any simple $C$-module $W$. In particular, $l\cdot W$ is not equal to 0 for any simple $C$-module $W$. By [16, Theorem 2.8.7], the lCllCl-$C$-bimodule lClC induces a Morita equivalence between $lCl$ and $C$. We are done. ∎
@@
By [32, Theorem 1], the hyperfocal subalgebra $\mathbb{A}$ is a relatively $\mbox{$\mathcal{O}$}D$-separable $D$-interior algebra. By [9, Lemma 3.2], $\mathbb{A}$ is an indecomposable $\mathcal{O}$-algebra. From the paragraph 3.1, we can get $D$ has a unique local point on $\mathbb{A}$. Since $A=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}u$, we have $\frac{\mathrm{rank}_{\mathcal{O}}(\mathbb{A})}{|D|}=\frac{\mathrm{rand}_{\mathcal{O}}(A)}{|P|}$ which is coprime to $p$ by [29, Corollary (44.8)]. Since $\mathbb{A}$ is a direct summand of $A$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules, the statements in the corollary follow from [17, Theorem 8.7.1] and Lemma 3.3 above. ∎
@@
Let us consider the quotient group $N_{G}(D_{\delta})/DC_{G}(D)$. Different from the inertial quotient $N_{G}(P_{\gamma})/PC_{G}(P)$ being a $p^{\prime}$-group, $N_{G}(D_{\delta})/DC_{G}(D)$ may have a nontrivial Sylow $p$-subgroup $PC_{G}(D)/DC_{G}(D)$ in general. However, we can show that it is a $p$-nilpotent group, which means that it has a normal $p$-complement. We call this normal $p$-complement a hyperfocal inertial quotient of the block $b$ and denote it by $E_{\mathfrak{h}}$ throughout this paper. Hence, $E_{\mathfrak{h}}$ is a $p^{\prime}$-group and $N_{G}(D_{\delta})/DC_{G}(D)=PC_{G}(D)/DC_{G}(D)\cdot E_{\mathfrak{h}}$. To show this result, we need the following lemma about the structures of the Brauer quotients $A_{\delta}(D)$ and $\mathbb{A}_{\delta}(D)$ of $A_{\delta}$ and $\mathbb{A}_{\delta}$ at $D$.
@@
Let $\bar{b}_{D}$ be the block of $C_{G}(D)$ over $k$ associated with the local pointed group $D_{\delta}$. It is clear that $\bar{b}_{D}$ is $P$-stable and nilpotent by [25, Proposition 3.3]. Then $\bar{b}_{D}$ can be viewed as a block of $PC_{G}(D)$ and we denote it by $\bar{c}_{D}$ when it is viewed as a block of $PC_{G}(D)$. So the block $\bar{c}_{D}$ is also nilpotent by [15, Proposition 6.5]. It is easy to check that the block $\bar{c}_{D}$ has a defect group $P$ and then $C_{P}(D)$ is a defect group of the block $\bar{b}_{D}$.
By [29, Lemma (40.2)], we have $\mathrm{Br}_{D}(i)$ belongs to a local point of $P$ on $kC_{G}(D)$. In particular, $\mathrm{Br}_{D}(i)kPC_{G}(D)\mathrm{Br}_{D}(i)$ is a source algebra of the block $\bar{c}_{D}$. Note that $A(D)=\mathrm{Br}_{D}(i)kC_{G}(D)\mathrm{Br}_{D}(i)$ is a primitive $P$-algebra. Then by [17, Theorem 8.12.3], there is a primitive Dade $P$-interior algebra $T_{D}$ over $k$ such that we have an isomorphism of $P$-interior algebras
$\mathrm{Br}_{D}(i)kPC_{G}(D)\mathrm{Br}_{D}(i)\cong T_{D}\mathop{\otimes}\limits_{k}kP$
@@
Since $D$ is normal in $PC_{G}(D)$, we have $v\cdot 1_{T_{D}}=1_{T_{D}}$ for any $v\in D$. By the isomorphism (3.6.1), we have a $Z(D)$-interior algebra isomorphism $A_{\delta}(D)\cong kC_{P}(D)$. Obviously, the image of $\mathbb{A}_{\delta}(D)$ under this isomorphism contains $kZ(D)$.
Borrowing the notation from the paragraph 3.1, for any $u\in P$, there is an element $a_{u,\delta}$ of $A_{\delta}^{*}$ such that ${{}^{a_{u,\delta}}}(v\cdot j)=({{}^{u}}v)\cdot j$ for any $v\in D$ and $A_{\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}_{\delta}a_{u,\delta}$. It is clear that $\mathbb{A}_{\delta}a_{u,\delta}$ is a $D$-stable $\mathcal{O}$-module for any $u\in P$ and then $A_{\delta}(D)=\mathop{\bigoplus}\limits_{uD\in P/D}(\mathbb{A}_{\delta}a_{u,\delta})(D)$. When $u$ is in $C_{P}(D)$, $a_{u,\delta}$ is in $A_{\delta}^{D}$ and then $(\mathbb{A}_{\delta}a_{u,\delta})(D)=\mathbb{A}_{\delta}(D)\mathrm{Br}_{D}(a_{u,\delta})$ and in particular they have the same dimension. Therefore, $\mathop{\bigoplus}\limits_{uZ(D)\in C_{P}(D)/Z(D)}(\mathbb{A}_{\delta}a_{u,\delta})(D)$ has dimension greater than or equal to $|C_{P}(D)|$. Then we are done. ∎
Now we can show that the quotient group $N_{G}(D_{\delta})/DC_{G}(D)$ is a $p$-nilpotent group.
@@
The quotient group $N_{G}(D_{\delta})/DC_{G}(D)$ is a $p$-nilpotent group.

### v2

@@ 3 Hyperfocal abelian Frobenius blocks
Let $G$ be a finite group and $b$ a block of $G$ over $\mathcal{O}$ with defect pointed group $P_{\gamma}$. We denote the hyperfocal subgroup of $P_{\gamma}$ by $D$. Then by [25, Propositions 4.2 and 3.3], there is a unique local point $\delta$ of $D$ on $\mbox{$\mathcal{O}$}Gb$ such that $D_{\delta}\leq P_{\gamma}$. Take an element $i$ of $\gamma$ and set $A=i\mbox{$\mathcal{O}$}Gi$. Then $A$ is a source algebra of $\mbox{$\mathcal{O}$}Gb$ and let $\mathbb{A}$ be a hyperfocal subalgebra of $A$. We take an element $j$ of $\delta$ such that $ij=j=ji$. By [25, Proposition 2.8], we can assume that $j$ belongs to $\mathbb{A}$. We can view $D_{\delta}$ as a local pointed group of $D$ on both $A$ and $\mathbb{A}$. Set $A_{\delta}=jAj$ and $\mathbb{A}_{\delta}=j\mathbb{A}j$. Clearly, $A_{\delta}$ and $\mathbb{A}_{\delta}$ are embedded algebras of the local pointed group $D_{\delta}$ on $A$ and $\mathbb{A}$, respectively. By the uniqueness of $D_{\delta}$, there is an element $a_{u}$ in $(\mathbb{A}^{D})^{\times}$ such that $a_{u}ja_{u}^{-1}=uju^{-1}$ for any $u\in P$. Denote $a_{u}^{-1}uj$ by $a_{u,\delta}$ for any $u\in P$. Then $a_{u,\delta}$ belongs to $A_{\delta}^{\times}$ fulfilling $a_{u,\delta}(vj)a_{u,\delta}^{-1}=(uvu^{-1})j$ for any $v\in D$ and $j\mathbb{A}uj=\mathbb{A}_{\delta}a_{u,\delta}$. Moreover, we have $A_{\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}_{\delta}a_{u,\delta}$ as $\mathcal{O}$-modules.
@@
It is well-known that the source algebra $A$ is closely related to the block algebra $\mbox{$\mathcal{O}$}Gb$. For example, they are both relatively $\mbox{$\mathcal{O}$}P$-separable $P$-interior algebras. There is a canonical Morita equivalence between the source algebra $A$ and the block algebra $\mbox{$\mathcal{O}$}Gb$ induced by the $A$-$\mbox{$\mathcal{O}$}Gb$-bimodule $i\mbox{$\mathcal{O}$}G$. Next, we will show that these properties also hold for the hyperfocal subalgebra $\mathbb{A}$ and its embedded subalgebra $\mathbb{A}_{\delta}$ by the following general result.
@@
Let $Q$ be a $p$-subgroup and $C$ an indecomposable relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebra. Assume that $C$ has a unique local pointed group $Q_{\epsilon}$ and as $\mbox{$\mathcal{O}$}(Q\times Q)$-modules, every indecomposable direct summand of $C$ has the form $\mathrm{Ind}_{\Delta_{\varphi}(Z)}^{Q\times Q}(\mbox{$\mathcal{O}$})$ for some subgroup $Z$ of $Q$ and some injective homomorphism $\varphi:Z\longrightarrow Q$. Here, $\Delta_{\varphi}(Z)=\{(u,\varphi(u))\mid u\in Z\}$ is a subgroup of $Q\times Q$. Moreover, suppose that $\frac{\mathrm{rank}_{\mathcal{O}}(C)}{|Q|}$ is coprime to $p$. Take an element $l$ of $\epsilon$. Then the $Q$-interior algebra $lCl$ is a relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebra and the lCllCl-$C$-bimodule lClC induces a Morita equivalence between $lCl$ and $C$.
@@
Since every indecomposable direct summand of the $\mbox{$\mathcal{O}$}(Q\times Q)$-module $C$ has the form $\mathrm{Ind}_{\Delta_{\varphi}(Z)}^{Q\times Q}(\mbox{$\mathcal{O}$})$, $C$ is a direct sum of some copies of $\mbox{$\mathcal{O}$}Q$ as left $\mbox{$\mathcal{O}$}Q$ and right $\mbox{$\mathcal{O}$}Q$-modules. In particular, $C$ is free as left and right $\mbox{$\mathcal{O}$}Q$-modules. Hence, the structural map $Q\longrightarrow C^{\times}$ is injective and we can view $\mbox{$\mathcal{O}$}Q$ as a subalgebra of $C$. Obviously, every finitely generated projective $C$-module has $\mathcal{O}$-rank divided by $|Q|$.
@@
By the definition of relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebras, $C$ is a direct summand of $C\mathop{\otimes}\limits_{\mathcal{O}Q}C$ as $C$-$C$-bimodules. Since $C$ is an indecomposable algebra, there are some primitive idempotents $e,f$ of $C^{Q}$ such that $C$ is a direct summand of $Ce\mathop{\otimes}\limits_{\mathcal{O}Q}fC$ as $C$-$C$-bimodules. Suppose that $\mathrm{Br}_{Q}(e)=0$. By [29, Theorem (23.1)], there is a proper subgroup $Q_{0}$ of $Q$ and a primitive idempotent $e_{0}$ of $C^{Q_{0}}$ such that $e=\mathrm{Tr}_{Q_{0}}^{Q}(e_{0})$ and ${{}^{u}}e_{0}\cdot e_{0}=0$ for any $u\in Q-Q_{0}$. Then the map sending $c\in Ce$ to $\sum\limits_{uQ_{0}\in Q/Q_{0}}cue_{0}\mathop{\otimes}\limits_{\mathcal{O}Q_{0}}u^{-1}$ induces a $C$-$\mbox{$\mathcal{O}$}Q$-bimodule isomorphism from $Ce$ to $Ce_{0}\mathop{\otimes}\limits_{\mathcal{O}Q_{0}}\mbox{$\mathcal{O}$}Q$. Therefore, $C$ is a direct summand of $Ce_{0}\mathop{\otimes}\limits_{\mathcal{O}Q_{0}}fC$ as $C$-$C$-bimodules. Denote by $\bar{C}$ the $k$-algebra $C/J(\mbox{$\mathcal{O}$})C$. Then for any simple $C$-module $W$, we have $W$ isomorphic to a direct summand of $\bar{C}\mathop{\otimes}\limits_{kQ_{0}}U$ for some indecomposable $kQ_{0}$-module $U$. By the assumption and Mackey’s formula, $\mathrm{Res}_{Q\times Q_{0}}^{Q\times Q}(\bar{C})$ is a direct sum of submodules isomorphic to ${\rm Ind}_{\Delta_{\psi}(Z)}^{Q\times Q_{0}}(k)$ for some proper subgroup $Z$ of $Q$ and injective homomorphism $\psi:Z\longrightarrow Q_{0}$. By Green’s indecomposability theorem, this implies that the dimension of every simple $C$-module has dimension divided by $p$. Let $P_{W}$ be the projective cover of the simple $C$-module $W$. We have $\mathrm{dim}(\bar{C})=\sum\limits_{W}\mathrm{dim}(P_{W})\mathrm{dim}(W)$, which is divided by $p|Q|$. This contradicts the assumption that $\frac{\mathrm{rank}_{\mathcal{O}}(C)}{|Q|}$ is coprime to $p$.
@@
We conclude that $\mathrm{Br}_{Q}(e)\neq 0$. Similarly, we can also get $\mathrm{Br}_{Q}(f)\neq 0$. Since $\epsilon$ is the unique local point of $Q$ on $C$, we have $e$ and $f$ both belonging to $\epsilon$. In particular, we have $Ce$ is isomorphic to $Cl$ as $C$-$\mbox{$\mathcal{O}$}Q$-bimodules and $fC$ is isomorphic to $lC$ as $\mbox{$\mathcal{O}$}Q$-$C$-bimodules. So $C$ is a direct summand of $Cl\mathop{\otimes}\limits_{\mathcal{O}Q}lC$ as $C$-$C$-bimodules and then $lCl$ is a direct summand of $lCl\mathop{\otimes}\limits_{\mathcal{O}Q}lCl$ as lCllCl-lCllCl-bimodules. This means that $lCl$ is a relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebra.
@@
Since $C$ is a direct summand of $Cl\mathop{\otimes}\limits_{\mathcal{O}Q}lC$, $W$ is a direct summand of $Cl\mathop{\otimes}\limits_{\mathcal{O}Q}l\cdot W$ for any simple $C$-module $W$. In particular, $l\cdot W$ is not equal to 0 for any simple $C$-module $W$. By [15, Theorem 2.8.7], the lCllCl-$C$-bimodule lClC induces a Morita equivalence between $lCl$ and $C$. ∎
@@
By [32, Theorem 1], the hyperfocal subalgebra $\mathbb{A}$ is a relatively $\mbox{$\mathcal{O}$}D$-separable $D$-interior algebra. By [9, Lemma 3.2], $\mathbb{A}$ is an indecomposable $\mathcal{O}$-algebra. From the paragraph 3.1, we can get that $D$ has a unique local point on $\mathbb{A}$. Since $A=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}u$, we have $\frac{\mathrm{rank}_{\mathcal{O}}(\mathbb{A})}{|D|}=\frac{\mathrm{rank}_{\mathcal{O}}(A)}{|P|}$ which is coprime to $p$ by [29, Corollary (44.8)]. Since $\mathbb{A}$ is a direct summand of $A$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules, the statements in the corollary follow from [16, Theorem 8.7.1] and Lemma 3.3 above. ∎
@@
Let us consider the quotient group $N_{G}(D_{\delta})/DC_{G}(D)$. Generally, $N_{G}(D_{\delta})/DC_{G}(D)$ may have a nontrivial Sylow $p$-subgroup $PC_{G}(D)/DC_{G}(D)$. This property is different from the fact that the inertial quotient $N_{G}(P_{\gamma})/PC_{G}(P)$ is a $p^{\prime}$-group. However, we can show that it is a $p$-nilpotent group, which means that it has a normal $p$-complement. We call this normal $p$-complement a hyperfocal inertial quotient of the block $b$ and denote it by $E_{\mathfrak{h}}$ throughout this paper. Hence, $E_{\mathfrak{h}}$ is a $p^{\prime}$-group and $N_{G}(D_{\delta})/DC_{G}(D)=PC_{G}(D)/DC_{G}(D)\cdot E_{\mathfrak{h}}$. To show this result, we need the following lemma about the structures of the Brauer quotients $A_{\delta}(D)$ and $\mathbb{A}_{\delta}(D)$ of $A_{\delta}$ and $\mathbb{A}_{\delta}$ at $D$, respectively.
@@
Let $\bar{b}_{D}$ be the block of $C_{G}(D)$ over $k$ associated with the local pointed group $D_{\delta}$. It is clear that $\bar{b}_{D}$ is $P$-stable and nilpotent by [25, Proposition 3.3]. Then $\bar{b}_{D}$ can be viewed as a block of $PC_{G}(D)$ and we denote it by $\bar{c}_{D}$ when it is viewed as a block of $PC_{G}(D)$. So the block $\bar{c}_{D}$ is also nilpotent by [14, Proposition 6.5]. It is easy to check that the block $\bar{c}_{D}$ has a defect group $P$ and then $C_{P}(D)$ is a defect group of the block $\bar{b}_{D}$.
By [29, Lemma (40.2)], we have $\mathrm{Br}_{D}(i)$ belongs to a local point of $P$ on $kC_{G}(D)$. In particular, $\mathrm{Br}_{D}(i)kPC_{G}(D)\mathrm{Br}_{D}(i)$ is a source algebra of the block $\bar{c}_{D}$. Note that $A(D)=\mathrm{Br}_{D}(i)kC_{G}(D)\mathrm{Br}_{D}(i)$ is a primitive $P$-algebra. Then by [16, Theorem 8.12.3], there is a primitive Dade $P$-interior algebra $T_{D}$ over $k$ such that we have an isomorphism of $P$-interior algebras
$\mathrm{Br}_{D}(i)kPC_{G}(D)\mathrm{Br}_{D}(i)\cong T_{D}\mathop{\otimes}\limits_{k}kP,$
@@
Since $D$ is normal in $PC_{G}(D)$, we have $v\cdot 1_{T_{D}}=1_{T_{D}}$ for any $v\in D$. By the isomorphism (3.6.1), we have a $Z(D)$-interior algebra isomorphism $A_{\delta}(D)\cong kC_{P}(D)$. Obviously, the image of $\mathbb{A}_{\delta}(D)$ under this isomorphism contains $kZ(D)$. In particular, the dimension ${\rm dim}_{k}(\mathbb{A}_{\delta}(D))$ of the $k$-algebra $\mathbb{A}_{\delta}(D)$ is greater than or equal to $|Z(D)|$.
Borrowing the notation from the paragraph 3.1, for any $u\in P$, there is an element $a_{u,\delta}$ in $A_{\delta}^{\times}$ such that ${{}^{a_{u,\delta}}}(v\cdot j)=({{}^{u}}v)\cdot j$ for any $v\in D$ and $A_{\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}_{\delta}a_{u,\delta}$. It is clear that $\mathbb{A}_{\delta}a_{u,\delta}$ is a $D$-stable $\mathcal{O}$-module for any $u\in P$ and then $A_{\delta}(D)=\mathop{\bigoplus}\limits_{uD\in P/D}(\mathbb{A}_{\delta}a_{u,\delta})(D)$. When $u$ is in $C_{P}(D)$, $a_{u,\delta}$ is in $(A_{\delta}^{D})^{\times}$ and then $(\mathbb{A}_{\delta}a_{u,\delta})(D)=\mathbb{A}_{\delta}(D)\mathrm{Br}_{D}(a_{u,\delta})$. In particular, they have the same dimension. Therefore, the $k$-algebra $A_{\delta}(D)=\mathop{\bigoplus}\limits_{uZ(D)\in C_{P}(D)/Z(D)}(\mathbb{A}_{\delta}a_{u,\delta})(D)$ has dimension greater than or equal to $|C_{P}(D)|$. Then we are done. ∎
Recall from the paragraph 2.12 that there is a group isomorphism $\mathfrak{f}_{G,D}$ from $N_{G}(D_{\delta})/DC_{G}(D)$ and $F_{\mathcal{O}G}(D)$. It is clear that $j\mbox{$\mathcal{O}$}Gj=jAj=A_{\delta}$. Then $F_{\mathcal{O}G}(D)=F_{A}(D)$ and $\mathfrak{f}_{G,D}$ is a group isomorphism from $N_{G}(D_{\delta})/DC_{G}(D)$ to $F_{A}(D)$. Now we can show that the quotient group $N_{G}(D_{\delta})/DC_{G}(D)$ is a $p$-nilpotent group.
@@
The quotient group $N_{G}(D_{\delta})/DC_{G}(D)$ is a $p$-nilpotent group. Moreover, the hyperfocal inertial quotient $E_{\mathfrak{h}}$ of the block $b$ is the preimage of the subgroup $N_{\mathbb{A}_{\delta}}(D)\cdot(A_{\delta}^{D})^{\times}/D(A_{\delta}^{D})^{\times}$ of $F_{A}(D)$ under the isomorphism $\mathfrak{f}_{G,D}$.